OpenStudy (anonymous):

A chemist has two solutions of hcl .One has a 40% concentration and the other a 25% concentration.How many litters must be mixed to obtain 136 liters of a 26% solution?

OpenStudy (anonymous):

can you do 2 equations and 2 unknowns?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. A and B are the amounts of the two solutions. A and B together form 136 liters of solution (eq #2). The first equation shows mixing an unknown amount A of 40% solution with an unknown amount B of 25% solution to get a known amount (136 liters) of 26% solution. Solve for A and B. Eq 1: 0.40A + 0.25B = 0.26(136) Eq 2: A + B = 136